The surface of the Sun has a temperature of about 5800 K. The radius of the Sun is 6.96 x 108 m. Calculate the total energy radiated by the Sun each second. Assume the emissivity of the Sun is 0.965

1
Expert's answer
2022-02-09T16:15:08-0500

Given:

"T=\rm 5800\:\rm K"

"R=\rm 6.96*10^8\:\rm m"

"\epsilon=0.965"

"\sigma=5.67*10^{-8}\:\rm W/(m^2\cdot K^4)"

"t=1\:\rm s"


The luminosity (radiation power) of Sun

"L=\epsilon \sigma T^4A=\epsilon \sigma T^4*4\pi R^2"

The total energy radiated by the Sun each second

"E=Lt=\epsilon \sigma T^4*4\pi R^2t"

"E=0.965*5.67*10^{-8}*5800^4\\
*4\pi*(6.96*10^8)^2*1=3.77*10^{26}\: \rm J"